A researcher is interested in testing the claim that more than 80% of the adults believe in global warming. She conducted a survey on a random sample of 800 adults. The survey showed that 650 adults in the sample believe in global warming. Assume the researcher wants to use a 0.05 significance level to test the claim.
(a)Identify the null hypothesis and the alternative hypothesis.
(b)Determine the test statistic.(Show all work)
(c)Determine the P-value for this test. (Show all work)
(d)Is there sufficient evidence to support the claim that more than 80% of adults believe in global
warming? Explain.

Answer by stanbon(75887)

A researcher is interested in testing the claim that more than 80% of the adults believe in global warming.
She conducted a survey on a random sample of 800 adults. The survey showed that 650 adults in the sample believe in global warming. Assume the researcher wants to use a 0.05 significance level to test the claim.
---
Sample proportion = 650/800 = 0.8125
Std = sqrt(0.8*0.2/800)= 0.0141
---------------------------
(a)Null hypothesis and the alternative hypothesis.
Ho: p = 0.8 (claim)
Ha:: p # 0.8
----------------------
(b)Determine the test statistic.(Show all work)
z(0.8125) = (0.8125-0.8)/0.0141 = 0.8865
-----
p-value = 2*p(z > 0.8865) = 2*normalcdf(0.8865,100) = 0.375
-----
Conclusion Since the p-value is greater than 5%, fail to reject Ho.
Conclusion:: The test results support the claim that p = 0.8
-----------------
Cheers,
Stan H.
------------

(c)Determine the P-value for this test. (Show all work)
(d)Is there sufficient evidence to support the claim that more than 80% of adults believe in global
warming? Explain.